Can someone please help me figure out how to work this out:
Find the 0.99 confidence interval for .
In an effort to compare college costs in State of Michigan, a sample of 36 junior students is randomly selected statewide from the private colleges and 36 more from the public colleges. The private college sample resulted in a mean of $27,650 and the public college sample mean was $11,360.Assume the annual college fees for private colleges have a mounded distribution and the standard deviation is $1725. Find the 95% confidence interval for the mean costs for private colleges.


Thanks!

Answer by stanbon(75887)

In an effort to compare college costs in State of Michigan, a sample of 36 junior students is randomly selected statewide from the private colleges and 36 more from the public colleges. The private college sample resulted in a mean of $27,650 and the public college sample mean was $11,360.Assume the annual college fees for private colleges have a mounded distribution and the standard deviation is $1725. Find the 95% confidence interval for the mean costs for private colleges.
----------------------------
sample mean:: 27,650
E = z*s/sqrt(n) = 1.96*1725/sqrt(36) = 563.5
-----
95%CI: 27,650-563.5 < u < 27650+563.5
------------------------------------
Cheers,
Stan H.
--------------------